OpenStudy (anonymous):

Find the value of the acute angle for which cos=sin(theta-20)

OpenStudy (anonymous):

I know the answer is 55 degrees, I just dont understand why!

OpenStudy (dumbcow):

its the identity \[\cos x = \sin (90-x)\] if the 2 angles add up to 90...then sin = cos therefore x + (x-20) = 90 --> x = 55

OpenStudy (anonymous):

thanks!

OpenStudy (dumbcow):

yw